my teacher gave me the expression .5(2x-14)=6x+28. i got the answer of x=-7...is that correct? and she told us to discuss the answer but i dont know how to do that.
Found 2 solutions by rothauserc, ewatrrr:
Answer by rothauserc(4718)

.5(2x-14) = 6x+28
multiply .5 times the items in parenthesis
x - 7 = 6x + 28
add 7 to both sides of =
x = 6x +35
subtract 6x from both sides of =
-5x = 35
divide both sides of = by -5
x = -7

Answer by ewatrrr(24785)

5(2x-14)=6x+28.
10x - 70 = 6x + 28
4x = 98
x = 98/4 = 49/2 = 24 1/2 = 24.5
The answer can be expressed as :
1) an Improper Fraction.
2) a Mixed Fraction
0r
3) as a Decimal.